# Temporal Smoothness-Aware Rate-Distortion Optimized 4D Gaussian Splatting

### v1

@@ 4.3 Ablation Study
Opacity Quantization. We validate the opacity quantization strategy described in Section 3.2 through ablation experiments summarized in Table 1. Specifically, we examine the quantization sensitivity of four parameters: static opacity ($\alpha^{s}$), dynamic opacity ($\alpha^{d}$), center parameters ($a_{s}^{o},a_{f}^{o}$), and variance parameters ($b_{s}^{o},b_{f}^{o}$). Although static and dynamic opacity share similar characteristics, we evaluate them separately, considering the potential heightened sensitivity of dynamic opacity. The results demonstrate that quantizing additional opacity-related parameters consistently improves storage efficiency. However, notably in the last two rows, quantizing variance parameters significantly reduces PSNR despite achieving marginal additional storage benefits. Based on the observation, we conclude that quantization should be limited to the static and dynamic opacity and center parameters to maintain a balanced trade-off between rendering fidelity and compression efficiency.
@@
Wavelet Transform for Dynamic Positions. We further evaluate the effectiveness of wavelet-based compression applied specifically to the dynamic position trajectories ($\boldsymbol{\mu}_{d}(t)$). As detailed in Section 3, our proposed method leverages a Haar wavelet transform to capture low-frequency motion components while explicitly discarding high-frequency details. Figure 3 compares the rate-distortion performance of our method with wavelet transform against the baseline without it, clearly highlighting the benefit of incorporating wavelet compression. Notably, employing the wavelet transform consistently improves performance across all tested compression levels, not only reducing the storage size but also enhancing rendering fidelity. Specifically, at compression level 1, our method achieves approximately 19% storage reduction alongside a PSNR gain of 0.19 dB. Even at the highest fidelity setting (level 6), we observe around 26% storage savings coupled with a PSNR improvement of 0.09 dB. These results emphasize that exploiting the inherent smoothness of point trajectories through wavelet-based compression not only effectively reduces redundancy in dynamic positions but also enhances trajectory modeling accuracy, leading to improved overall RD performance.
@@ 4.4 Comparative Evaluation
Table 2: Quantitative results on the N3V dataset comparing PSNR (dB), average model size (MB), and FPS. Our proposed RD-optimized method offers flexible compression levels. Level 6 achieves a significant size reduction (e.g., about 10$\times$ reduction to 11.06MB) compared to the Ex4DGS ex4dgs while maintaining competitive quality. Level 1 demonstrates the capability for aggressive compression (about 90$\times$ size reduction to 1.26MB) for highly resource-constrained scenarios.
@@
Ours [$\mathbb{L}$1] 24.97 28.38 28.02 24.16 28.33 28.39 27.04 1.26 (-98.9%) 163.0
Ours [$\mathbb{L}$6] 26.17 30.69 31.45 26.58 31.34 31.70 29.66 11.06 (-90.4%) 100.9
@@
We report the PSNR, storage size, and FPS for the N3V dataset in Table 2. The most notable advantage of our method is its significant size reduction. At compression Level 1, our model achieves a remarkable 98.9% size reduction compared to the baseline Ex4DGS, resulting in an average size of only 1.26 MB. Even when compared to 4DGaussians, the previously most storage-efficient method, our approach demonstrates over 25 times greater efficiency at the same compression level. Our method significantly outperforms existing models in rendering speed, reaching up to 163 FPS at Level 1, suitable for real-time applications.
Table 3: Quantitative results averaged over the Technicolor dataset. We compare the proposed method (L1, L6) against the Ex4DGS ex4dgs using standard quality metrics and model size.
@@
Ours [$\mathbb{L}$6] 32.20 0.904 0.100 19.6(-86.0%) 113.1
Ours [$\mathbb{L}$1] 28.60 0.822 0.232 2.1(-98.5%) 213.9
While there is an expected trade-off in rendering quality (PSNR) due to aggressive compression, our model still produces results that are sufficiently competitive. Particularly noteworthy is the comparison at Level 6, where our method not only has a smaller model size (11.06 MB) but also a higher average PSNR (29.66 dB) compared to 4DGaussians (34 MB, 28.63 dB).
@@ 5 Discussion
Conclusion. We presented the first end-to-end rate-distortion (RD) optimized compression framework for 4D Gaussian Splatting (4DGS) to address the challenge of large model sizes hindering deployment. Leveraging Ex4DGS, our method uses static/dynamic decomposition with adaptive RD-guided quantization and pointwise wavelet transform to achieve flexible compression. Our approach demonstrates significant model size reduction (down to approximately 1.1% of the original size ) while maintaining reasonable visual quality across standard benchmarks and offering user control over the rate-quality trade-off. This significantly improves the practicality of 4DGS for real-time rendering on diverse computational platforms, including resource-constrained devices.
Limitation. We observed motion blur artifacts of dynamic objects at high compression levels. Additionally, dynamic points still account for a substantial portion of total storage, highlighting significant room for improvement in compressing dynamic components (Appendices D and E). Future work includes refining the loss function, for example by applying separate masking weights for static and dynamic points, to further enhance dynamic point compression and overall visual quality.
@@
Appendix A Additional Quantitative Evaluation Results
We report the detailed quantitative evaluation results, including storage size and rendering fidelity metrics (PSNR, SSIM, LPIPS), for each scene of the N3V and Technicolor datasets in Table 4. The most notable highlight is the substantial model size reduction achieved by our proposed method across all scenes, ranging from 79.3% at the highest fidelity setting (Level 6) to 99.3% at the most aggressive compression level (Level 1). Although aggressive compression (lower levels) occasionally leads to significant drops in rendering fidelity, our framework provides multiple trade-off options to flexibly adapt according to deployment scenarios and resource constraints. These results indicate promising directions for future work, focusing on enhancing reconstruction fidelity at lower compression levels to further broaden applicability.
@@ Appendix A Additional Quantitative Evaluation Results
Ours [$\mathbb{L}$6] 0.897 26.17 0.091 15.29 (-88.2%) Ours [$\mathbb{L}$6] 0.942 30.69 0.064 10.44 (-91.2%)
Ours [$\mathbb{L}$5] 0.894 25.92 0.098 10.60 (-91.8%) Ours [$\mathbb{L}$5] 0.939 30.50 0.069 6.75 (-94.3%)
Ours [$\mathbb{L}$4] 0.892 25.82 0.101 7.40 (-94.3%) Ours [$\mathbb{L}$4] 0.935 30.11 0.075 4.50 (-96.2%)
Ours [$\mathbb{L}$3] 0.888 25.74 0.108 5.21 (-96.0%) Ours [$\mathbb{L}$3] 0.930 29.55 0.083 3.12 (-97.4%)
Ours [$\mathbb{L}$2] 0.882 25.48 0.119 3.48 (-97.3%) Ours [$\mathbb{L}$2] 0.922 28.99 0.099 2.06 (-98.3%)
Ours [$\mathbb{L}$1] 0.868 24.97 0.144 1.92 (-98.5%) Ours [$\mathbb{L}$1] 0.910 28.38 0.121 1.15 (-99.0%)
@@
Ours [$\mathbb{L}$6] 0.943 31.45 0.064 10.04 (-91.8%) Ours [$\mathbb{L}$6] 0.910 26.58 0.079 13.43 (-89.5%)
Ours [$\mathbb{L}$5] 0.940 30.86 0.068 6.59 (-94.6%) Ours [$\mathbb{L}$5] 0.908 26.35 0.082 8.97 (-93.0%)
Ours [$\mathbb{L}$4] 0.936 30.23 0.074 4.42 (-96.4%) Ours [$\mathbb{L}$4] 0.905 25.92 0.088 6.07 (-95.3%)
Ours [$\mathbb{L}$3] 0.931 29.82 0.082 3.03 (-97.5%) Ours [$\mathbb{L}$3] 0.902 25.68 0.095 4.18 (-96.7%)
Ours [$\mathbb{L}$2] 0.924 29.04 0.093 2.01 (-98.4%) Ours [$\mathbb{L}$2] 0.893 24.87 0.107 2.77 (-97.8%)
Ours [$\mathbb{L}$1] 0.911 28.02 0.122 1.12 (-99.1%) Ours [$\mathbb{L}$1] 0.875 24.16 0.138 1.52 (-98.8%)
@@
Ours [$\mathbb{L}$6] 0.950 31.34 0.059 8.49 (-91.6%) Ours [$\mathbb{L}$6] 0.950 31.70 0.056 8.66 (-90.8%)
Ours [$\mathbb{L}$5] 0.947 30.86 0.063 5.41 (-94.7%) Ours [$\mathbb{L}$5] 0.948 31.22 0.061 5.43 (-94.2%)
Ours [$\mathbb{L}$4] 0.943 30.38 0.070 3.56 (-96.5%) Ours [$\mathbb{L}$4] 0.944 30.62 0.068 3.58 (-96.2%)
Ours [$\mathbb{L}$3] 0.938 29.95 0.080 2.40 (-97.6%) Ours [$\mathbb{L}$3] 0.938 29.95 0.077 2.48 (-97.3%)
Ours [$\mathbb{L}$2] 0.931 29.13 0.093 1.58 (-98.4%) Ours [$\mathbb{L}$2] 0.931 29.34 0.093 1.64 (-98.3%)
Ours [$\mathbb{L}$1] 0.918 28.33 0.122 0.89 (-99.1%) Ours [$\mathbb{L}$1] 0.917 28.39 0.122 0.94 (-99.0%)
@@
Ours [$\mathbb{L}$6] 0.931 31.55 0.040 33.19 (-84.6%)
Ours [$\mathbb{L}$5] 0.928 31.38 0.045 21.76 (-89.9%)
Ours [$\mathbb{L}$4] 0.913 30.39 0.060 12.73 (-94.1%)
Ours [$\mathbb{L}$3] 0.889 29.86 0.092 7.61 (-96.5%)
Ours [$\mathbb{L}$2] 0.845 28.44 0.146 4.55 (-97.9%)
Ours [$\mathbb{L}$1] 0.765 26.32 0.247 2.32 (-98.9%)
@@
Ours [$\mathbb{L}$6] 0.932 31.33 0.046 33.62 (-79.3%) Ours [$\mathbb{L}$6] 0.871 32.90 0.168 5.89 (-92.9%)
Ours [$\mathbb{L}$5] 0.932 31.33 0.047 25.47 (-84.3%) Ours [$\mathbb{L}$5] 0.867 32.74 0.182 2.86 (-96.6%)
Ours [$\mathbb{L}$4] 0.929 31.08 0.051 18.50 (-88.6%) Ours [$\mathbb{L}$4] 0.861 32.60 0.200 1.77 (-97.9%)
Ours [$\mathbb{L}$3] 0.924 31.00 0.057 13.59 (-91.6%) Ours [$\mathbb{L}$3] 0.856 32.06 0.214 1.27 (-98.5%)
Ours [$\mathbb{L}$2] 0.916 30.56 0.068 9.34 (-94.2%) Ours [$\mathbb{L}$2] 0.850 31.68 0.228 0.90 (-98.9%)
Ours [$\mathbb{L}$1] 0.896 29.46 0.092 5.39 (-96.7%) Ours [$\mathbb{L}$1] 0.840 30.44 0.254 0.58 (-99.3%)
@@
Ours [$\mathbb{L}$6] 0.914 35.05 0.110 9.19 (-91.4%) Ours [$\mathbb{L}$6] 0.871 30.16 0.138 16.27 (-87.8%)
Ours [$\mathbb{L}$5] 0.905 34.00 0.122 5.55 (-94.8%) Ours [$\mathbb{L}$5] 0.869 30.12 0.150 9.31 (-93.0%)
Ours [$\mathbb{L}$4] 0.891 33.29 0.142 3.47 (-96.7%) Ours [$\mathbb{L}$4] 0.858 29.70 0.172 5.58 (-95.8%)
Ours [$\mathbb{L}$3] 0.873 32.20 0.168 2.30 (-97.8%) Ours [$\mathbb{L}$3] 0.849 29.42 0.191 3.57 (-97.3%)
Ours [$\mathbb{L}$2] 0.848 30.96 0.210 1.47 (-98.6%) Ours [$\mathbb{L}$2] 0.833 29.00 0.221 2.19 (-98.4%)
Ours [$\mathbb{L}$1] 0.804 28.93 0.293 0.79 (-99.3%) Ours [$\mathbb{L}$1] 0.804 27.84 0.274 1.16 (-99.1%)
Appendix B Rate-Distortion Performance for Each Scene
In addition to the numerical results provided in Appendix A, we illustrate detailed per-scene rate-distortion (RD) curves for both datasets. Specifically, Figures 6 and 7 show RD curves for the N3V and Technicolor datasets, respectively. These figures visually reaffirm the notable size reductions achieved by our method, consistently demonstrating drastic reductions in model sizes. A closer examination reveals that fidelity degradation for the Technicolor dataset is relatively moderate compared to N3V, likely due to the simpler and less intricate motions inherent in Technicolor scenes. Conversely, N3V contains complex and detailed actions, such as precise tool movements or dynamic interactions (e.g., operating kitchen tongs or using a torch to sear meat), making compression-induced fidelity degradation more pronounced. This observation highlights opportunities for further enhancing compression methods, specifically targeting more effective exploitation of detailed volumetric motion information. Another noteworthy finding appears in the Technicolor dataset’s “Train” scene, where certain compression levels achieve superior rendering fidelity compared to the baseline Ex4DGS method. This result indicates the potential benefit of model simplicity, characterized by fewer points and lower entropy, which can unexpectedly enhance view-synthesis quality by effectively reducing model redundancy.

### v2

@@ 4.3 Ablation Study
Opacity Quantization. We validate the opacity quantization strategy described in Section˜3.2 through ablation experiments summarized in Table˜1. Recognizing that each parameter contributes differently to the final visual quality, we designed this experiment to identify the optimal subset of parameters for quantization that best balances compression efficiency and rendering fidelity, rather than naively quantizing all components. Specifically, we examine the quantization sensitivity of four parameters: static opacity ($\alpha^{s}$), dynamic opacity ($\alpha^{d}$), center parameters ($a_{s}^{o},a_{f}^{o}$), and variance parameters ($b_{s}^{o},b_{f}^{o}$). Although static and dynamic opacity share similar characteristics, we evaluate them separately, considering the potential heightened sensitivity of dynamic opacity. The results demonstrate that quantizing additional opacity-related parameters consistently improves storage efficiency. However, notably in the last two rows, quantizing variance parameters significantly reduces PSNR despite achieving marginal additional storage benefits. Based on the observation, we conclude that quantization should be limited to the static and dynamic opacity and center parameters to maintain a balanced trade-off between rendering fidelity and compression efficiency.
@@
Wavelet Transform for Dynamic Positions. We further evaluate the effectiveness of wavelet-based compression applied specifically to the dynamic position trajectories ($\boldsymbol{\mu}_{d}(t)$). As detailed in Section˜3, our proposed method leverages a Haar wavelet transform to capture low-frequency motion components while explicitly discarding high-frequency details. Figure˜3 compares the rate-distortion performance of our method with wavelet transform against the baseline without it, clearly highlighting the benefit of incorporating wavelet compression. As shown in the figure, the rate-distortion curve of our method (’Ours’) is consistently positioned to the top-left of the baseline (’w/o Wavelet Transform’), indicating a superior performance trade-off. This positioning signifies that for any given model size, our wavelet-based approach yields a higher PSNR, and conversely, for any target PSNR, it requires a smaller model size. Notably, employing the wavelet transform consistently improves performance across all tested compression levels, not only reducing the storage size but also enhancing rendering fidelity. Specifically, at compression level 1, our method achieves approximately 19% storage reduction alongside a PSNR gain of 0.19 dB. Even at the highest fidelity setting (level 6), we observe around 26% storage savings coupled with a PSNR improvement of 0.09 dB. These results emphasize that exploiting the inherent smoothness of point trajectories through wavelet-based compression not only effectively reduces redundancy in dynamic positions but also enhances trajectory modeling accuracy, leading to improved overall RD performance.
Multi-Level Wavelet Decomposition. To further analyze the impact of our wavelet compression, we conduct an ablation study on the level of wavelet decomposition applied to dynamic position trajectories. In addition to our proposed single-level Haar wavelet transform (retaining 1/2 of the coefficients),
Table 2: Ablation study on multi-level wavelet decomposition for dynamic positions, averaged over the N3V dataset. Columns compare decompositions retaining different fractions of wavelet coefficients.
Compression Level 1. (1/2) Level 2. (1/4) Level 3. (1/8)
Level PSNR / Size PSNR / Size PSNR / Size
$\mathbb{L}$6 28.73 / 14.05 28.47 / 11.35 27.74 / 9.80
$\mathbb{L}$5 29.59 / 9.14 28.57 / 7.51 27.72 / 6.50
$\mathbb{L}$4 29.14 / 6.23 28.14 / 5.14 27.44 / 4.44
$\mathbb{L}$3 28.59 / 4.37 27.81 / 3.61 27.38 / 3.10
$\mathbb{L}$2 28.27 / 2.91 27.57 / 2.42 27.17 / 2.05
$\mathbb{L}$1 27.39 / 1.63 26.89 / 1.36 26.43 / 1.16
we test more aggressive 2-level (retaining 1/4) and 3-level (retaining 1/8) decompositions. As summarized in Table˜2, applying a deeper multi-level decomposition progressively reduces the model size. For instance, at compression Level 1, the size decreases from 1.63 MB to 1.16 MB. However, this gain in compression comes at a consistent cost to rendering quality. Crucially, the new operating points generated by the multi-level transforms do not surpass the RD curve established by our original single-level method.
Analysis on Higher-Fidelity Trade-offs. Our framework’s compression is achieved through multiple lossy mechanisms, including pruning, ECVQ, and the wavelet transform. Therefore, we explore whether adjusting parameters beyond pruning can yield higher-quality results. Specifically, starting from our Level 6 setting, we test two strategies: (1) reducing the rate-loss weight, $\lambda_{R}$, in our objective function, and (2) additionally increasing the ECVQ codebook size.
Table 3: Analysis of the RD trade-off at higher fidelities. We compare L6 configuration against variants with a reduced ($\lambda_{\text{R}}$) and an increased codebook size. Results are averaged over the N3V dataset. [^2] The PSNR and Size values for our Level 1 and 6 models differ from those in Table˜4 as a different pre-trained model was used for this ablation study. For a detailed explanation, please see Appendix˜F.
Configuration Avg PSNR (dB) Size (MB)
Ex4DGS (Baseline) 32.11 115.63
Ours ($\mathbb{L}$6, low $\lambda_{\text{R}}$, large codebook) 30.29 27.21
Ours ($\mathbb{L}$6, low $\lambda_{\text{R}}$) 30.04 14.40
Ours ($\mathbb{L}$6) 28.73 14.05
Ours ($\mathbb{L}$1) 27.39 1.63
The results, averaged on the N3V dataset, are presented in Table˜3. Reducing $\lambda_{R}$ significantly boosts the PSNR by 1.31 dB (from 28.73 to 30.04 dB) with only a marginal increase in model size (0.35 MB). While also enlarging the codebook provides a minor additional PSNR gain, it nearly doubles the model size, indicating a less efficient trade-off. This study demonstrates that our framework can be flexibly tuned to bridge the quality gap with the baseline by adjusting the rate-loss weight.
@@ 4.4 Comparative Evaluation
[…]
@@
Ours $\mathbb{L}$1] 24.97 28.38 28.02 24.16 28.33 28.39 27.04 1.26 (-98.9%) 163.0
Ours $\mathbb{L}$6] 26.17 30.69 31.45 26.58 31.34 31.70 29.66 11.06 (-90.4%) 100.9
@@
We report the PSNR, storage size, and FPS for the N3V dataset in Table˜4. The most notable advantage of our method is its significant size reduction. At compression Level 1, our model achieves a remarkable 98.9% size reduction compared to the baseline Ex4DGS, resulting in an average size of only 1.26 MB. Even when compared to 4DGaussians, the previously most storage-efficient method, our approach demonstrates over 25 times greater efficiency at the same compression level. Our method significantly outperforms existing models in rendering speed, reaching up to 163 FPS at Level 1, suitable for real-time applications. While there is an expected trade-off in rendering quality (PSNR) due to aggressive compression,
Table 5: Quantitative results averaged over the Technicolor dataset. We compare the proposed method ($\mathbb{L}$1, $\mathbb{L}$6) against the Ex4DGS ex4dgs using standard quality metrics and model size.
@@
Ours $\mathbb{L}$6] 32.20 0.904 0.100 19.6(-86.0%) 113.1
Ours $\mathbb{L}$1] 28.60 0.822 0.232 2.1(-98.5%) 213.9
our model still produces results that are sufficiently competitive. Particularly noteworthy is the comparison at Level 6, where our method not only has a smaller model size (11.06 MB) but also a higher average PSNR (29.66 dB) compared to 4DGaussians (34 MB, 28.63 dB). A similar trend is observed for the Technicolor dataset, as summarized in Table˜5.
@@ 5 Discussion
Conclusion. We presented the first end-to-end rate-distortion (RD) optimized compression framework for 4D Gaussian Splatting (4DGS) to address the challenge of large model sizes hindering deployment. Leveraging Ex4DGS, our method uses static/dynamic decomposition with adaptive RD-guided quantization and pointwise wavelet transform to achieve flexible compression. Although this work builds upon the Ex4DGS framework, the core principles of our rate-distortion optimized compression are general and can be adapted to other 4DGS models. Our approach demonstrates significant model size reduction (down to approximately 1.1% of the original size ) while maintaining reasonable visual quality across standard benchmarks and offering user control over the rate-quality trade-off. This significantly improves the practicality of 4DGS for real-time rendering on diverse computational platforms, including resource-constrained devices.
Limitation. Our approach of filtering high-frequency motion via a wavelet transform has an inherent limitation in handling excessively dynamic movements, which can manifest as motion blur artifacts on fast-moving objects, particularly at high compression levels. Additionally, dynamic points still account for a substantial portion of total storage, highlighting significant room for improvement in compressing dynamic components (Appendices˜G and H). Future work includes refining the loss function, for example by applying separate masking weights for static and dynamic points, to further enhance dynamic point compression and overall visual quality.
@@
Appendix A Detailed Rate Loss Formulation
In Section˜3.2, we introduced the rate term $\mathcal{L}_{\text{rate}}$, which is a core component of our end-to-end rate-distortion optimization framework. The structure of our rate loss is based on the formulation proposed for static scenes in wang2024rdo3dgs , which we adapt to the unique challenges of dynamic 4D Gaussian Splatting by extending it to handle both static and dynamic components separately.
This section provides the detailed formulation of $\mathcal{L}_{\text{rate}}$. The total rate loss is a weighted sum of four distinct terms: Gaussian pruning loss ($\mathcal{L}_{\text{GSprune}}$), SH coefficient pruning loss ($\mathcal{L}_{\text{SHprune}}$), an entropy loss ($\mathcal{L}_{\text{entropy}}$), and a vector quantization loss ($\mathcal{L}_{\text{VQ}}$).
The complete expression for the total rate loss is:
$\mathcal{L}_{\text{rate}}=\lambda_{\text{GSprune}}\mathcal{L}_{\text{GSprune}}+\lambda_{\text{SHprune}}\mathcal{L}_{\text{SHprune}}+\mathcal{L}_{\text{entropy}}+\mathcal{L}_{\text{VQ}}$ (10)
Here, $\mathcal{L}_{\text{GSprune}}$ and $\mathcal{L}_{\text{SHprune}}$ are the pruning losses defined in Equations (4) and (5) of the main text, which encourage model sparsity. The terms $\mathcal{L}_{\text{entropy}}$ and $\mathcal{L}_{\text{VQ}}$ arise from the entropy-constrained vector quantization (ECVQ) applied to various Gaussian attributes (e.g., scale, rotation, SH coefficients).
The entropy loss, $\mathcal{L}_{\text{entropy}}$, estimates the bitrate required to encode the quantized parameters. It is formulated as the cross-entropy between the distribution of the quantized symbols and our learned entropy model:
$\small\mathcal{L}_{\text{entropy}}=\frac{1}{N}\sum_{m\in\{\text{static, dynamic}\}}\sum_{i=1}^{N}\left(\frac{r_{i,j}^{(s,m)}}{\lambda^{(s,m)}}+\frac{r_{i,j}^{(r,m)}}{\lambda^{(r,m)}}+\frac{r_{i,j}^{(DC,m)}}{\lambda^{(DC,m)}}+\frac{r_{i,j}^{(SH1,m)}}{\lambda^{(SH1,m)}}+\frac{r_{i,j}^{(SH2,m)}}{\lambda^{(SH2,m)}}+\frac{r_{i,j}^{(SH3,m)}}{\lambda^{(SH3,m)}}\right)$ (11)
The vector quantization loss, $\mathcal{L}_{\text{VQ}}$, measures the distortion introduced by mapping continuous parameter values to discrete codebook vectors during quantization:
$\mathcal{L}_{\text{VQ}}=\frac{1}{N}\sum_{m\in\{\text{static, dynamic}\}}\sum_{i=1}^{N}\left(d_{i,j}^{(s,m)}+d_{i,j}^{(r,m)}+d_{i,j}^{(DC,m)}+d_{i,j}^{(SH1,m)}+d_{i,j}^{(SH2,m)}+d_{i,j}^{(SH3,m)}\right)$ (12)
In these equations, the superscripts $(s,m),(r,m),\dots$ denote different attributes such as scale and rotation for both static and dynamic ($m$) Gaussians. The term $r_{i,j}$ represents the estimated bit cost (rate), while $d_{i,j}$ signifies the quantization error (distortion).
Appendix B Additional Quantitative Evaluation Results
We report the detailed quantitative evaluation results, including storage size and rendering fidelity metrics (PSNR, SSIM, LPIPS), for each scene of the N3V and Technicolor datasets in Table˜6. The most notable highlight is the substantial model size reduction achieved by our proposed method across all scenes, ranging from 79.3% at the highest fidelity setting (Level 6) to 99.3% at the most aggressive compression level (Level 1). Although aggressive compression (lower levels) occasionally leads to significant drops in rendering fidelity, our framework provides multiple trade-off options to flexibly adapt according to deployment scenarios and resource constraints. These results indicate promising directions for future work, focusing on enhancing reconstruction fidelity at lower compression levels to further broaden applicability.
@@ Appendix B Additional Quantitative Evaluation Results
Ours $\mathbb{L}$6] 0.897 26.17 0.091 15.29 (-88.2%) Ours $\mathbb{L}$6] 0.942 30.69 0.064 10.44 (-91.2%)
Ours $\mathbb{L}$5] 0.894 25.92 0.098 10.60 (-91.8%) Ours $\mathbb{L}$5] 0.939 30.50 0.069 6.75 (-94.3%)
Ours $\mathbb{L}$4] 0.892 25.82 0.101 7.40 (-94.3%) Ours $\mathbb{L}$4] 0.935 30.11 0.075 4.50 (-96.2%)
Ours $\mathbb{L}$3] 0.888 25.74 0.108 5.21 (-96.0%) Ours $\mathbb{L}$3] 0.930 29.55 0.083 3.12 (-97.4%)
Ours $\mathbb{L}$2] 0.882 25.48 0.119 3.48 (-97.3%) Ours $\mathbb{L}$2] 0.922 28.99 0.099 2.06 (-98.3%)
Ours $\mathbb{L}$1] 0.868 24.97 0.144 1.92 (-98.5%) Ours $\mathbb{L}$1] 0.910 28.38 0.121 1.15 (-99.0%)
@@
Ours $\mathbb{L}$6] 0.943 31.45 0.064 10.04 (-91.8%) Ours $\mathbb{L}$6] 0.910 26.58 0.079 13.43 (-89.5%)
Ours $\mathbb{L}$5] 0.940 30.86 0.068 6.59 (-94.6%) Ours $\mathbb{L}$5] 0.908 26.35 0.082 8.97 (-93.0%)
Ours $\mathbb{L}$4] 0.936 30.23 0.074 4.42 (-96.4%) Ours $\mathbb{L}$4] 0.905 25.92 0.088 6.07 (-95.3%)
Ours $\mathbb{L}$3] 0.931 29.82 0.082 3.03 (-97.5%) Ours $\mathbb{L}$3] 0.902 25.68 0.095 4.18 (-96.7%)
Ours $\mathbb{L}$2] 0.924 29.04 0.093 2.01 (-98.4%) Ours $\mathbb{L}$2] 0.893 24.87 0.107 2.77 (-97.8%)
Ours $\mathbb{L}$1] 0.911 28.02 0.122 1.12 (-99.1%) Ours $\mathbb{L}$1] 0.875 24.16 0.138 1.52 (-98.8%)
@@
Ours $\mathbb{L}$6] 0.950 31.34 0.059 8.49 (-91.6%) Ours $\mathbb{L}$6] 0.950 31.70 0.056 8.66 (-90.8%)
Ours $\mathbb{L}$5] 0.947 30.86 0.063 5.41 (-94.7%) Ours $\mathbb{L}$5] 0.948 31.22 0.061 5.43 (-94.2%)
Ours $\mathbb{L}$4] 0.943 30.38 0.070 3.56 (-96.5%) Ours $\mathbb{L}$4] 0.944 30.62 0.068 3.58 (-96.2%)
Ours $\mathbb{L}$3] 0.938 29.95 0.080 2.40 (-97.6%) Ours $\mathbb{L}$3] 0.938 29.95 0.077 2.48 (-97.3%)
Ours $\mathbb{L}$2] 0.931 29.13 0.093 1.58 (-98.4%) Ours $\mathbb{L}$2] 0.931 29.34 0.093 1.64 (-98.3%)
Ours $\mathbb{L}$1] 0.918 28.33 0.122 0.89 (-99.1%) Ours $\mathbb{L}$1] 0.917 28.39 0.122 0.94 (-99.0%)
@@
Ours $\mathbb{L}$6] 0.931 31.55 0.040 33.19 (-84.6%)
Ours $\mathbb{L}$5] 0.928 31.38 0.045 21.76 (-89.9%)
Ours $\mathbb{L}$4] 0.913 30.39 0.060 12.73 (-94.1%)
Ours $\mathbb{L}$3] 0.889 29.86 0.092 7.61 (-96.5%)
Ours $\mathbb{L}$2] 0.845 28.44 0.146 4.55 (-97.9%)
Ours $\mathbb{L}$1] 0.765 26.32 0.247 2.32 (-98.9%)
@@
Ours $\mathbb{L}$6] 0.932 31.33 0.046 33.62 (-79.3%) Ours $\mathbb{L}$6] 0.871 32.90 0.168 5.89 (-92.9%)
Ours $\mathbb{L}$5] 0.932 31.33 0.047 25.47 (-84.3%) Ours $\mathbb{L}$5] 0.867 32.74 0.182 2.86 (-96.6%)
Ours $\mathbb{L}$4] 0.929 31.08 0.051 18.50 (-88.6%) Ours $\mathbb{L}$4] 0.861 32.60 0.200 1.77 (-97.9%)
Ours $\mathbb{L}$3] 0.924 31.00 0.057 13.59 (-91.6%) Ours $\mathbb{L}$3] 0.856 32.06 0.214 1.27 (-98.5%)
Ours $\mathbb{L}$2] 0.916 30.56 0.068 9.34 (-94.2%) Ours $\mathbb{L}$2] 0.850 31.68 0.228 0.90 (-98.9%)
Ours $\mathbb{L}$1] 0.896 29.46 0.092 5.39 (-96.7%) Ours $\mathbb{L}$1] 0.840 30.44 0.254 0.58 (-99.3%)
@@
Ours $\mathbb{L}$6] 0.914 35.05 0.110 9.19 (-91.4%) Ours $\mathbb{L}$6] 0.871 30.16 0.138 16.27 (-87.8%)
Ours $\mathbb{L}$5] 0.905 34.00 0.122 5.55 (-94.8%) Ours $\mathbb{L}$5] 0.869 30.12 0.150 9.31 (-93.0%)
Ours $\mathbb{L}$4] 0.891 33.29 0.142 3.47 (-96.7%) Ours $\mathbb{L}$4] 0.858 29.70 0.172 5.58 (-95.8%)
Ours $\mathbb{L}$3] 0.873 32.20 0.168 2.30 (-97.8%) Ours $\mathbb{L}$3] 0.849 29.42 0.191 3.57 (-97.3%)
Ours $\mathbb{L}$2] 0.848 30.96 0.210 1.47 (-98.6%) Ours $\mathbb{L}$2] 0.833 29.00 0.221 2.19 (-98.4%)
Ours $\mathbb{L}$1] 0.804 28.93 0.293 0.79 (-99.3%) Ours $\mathbb{L}$1] 0.804 27.84 0.274 1.16 (-99.1%)
Appendix C Rate-Distortion Performance for Each Scene
In addition to the numerical results provided in Appendix˜B, we illustrate detailed per-scene rate-distortion (RD) curves for both datasets. Specifically, Figures˜6 and 7 show RD curves for the N3V and Technicolor datasets, respectively. These figures visually reaffirm the notable size reductions achieved by our method, consistently demonstrating drastic reductions in model sizes. A closer examination reveals that fidelity degradation for the Technicolor dataset is relatively moderate compared to N3V, likely due to the simpler and less intricate motions inherent in Technicolor scenes. Conversely, N3V contains complex and detailed actions, such as precise tool movements or dynamic interactions (e.g., operating kitchen tongs or using a torch to sear meat), making compression-induced fidelity degradation more pronounced. This observation highlights opportunities for further enhancing compression methods, specifically targeting more effective exploitation of detailed volumetric motion information. Another noteworthy finding appears in the Technicolor dataset’s “Train” scene, where certain compression levels achieve superior rendering fidelity compared to the baseline Ex4DGS method. This result indicates the potential benefit of model simplicity, characterized by fewer points and lower entropy, which can unexpectedly enhance view-synthesis quality by effectively reducing model redundancy.